If a right triangle has sides of 12 and 5, what is the hypotenuse side equal to?
If a right triangle has a hypotenuse side equal to 13 and one side equal to 5 what is the other side?


Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

If a right triangle has sides of 12 and 5, what is the hypotenuse side equal to?
h = sqrt(12^2+5^2) = sqrt(169) = 13
----------------------------------------------
If a right triangle has a hypotenuse side equal to 13 and one side equal to 5 what is the other side?
13^2 = x^2 + 5^2
x^2 = 169-25
x^2 = 144
x = 12


=======================================================
Cheers,
Stan H.

Answer by richwmiller(17219)

Your two questions really are one question.
one of the special right triangles you should know is the 5, 12, 13 right triangle.
http://www.friesian.com/pythag.htm